A negatively charged particle travels in a uniformly circular trajectory around an infinitely long positively charged rod with a linear charge density of lambda. The particle's speed (v) and distance (r) from the rod are constant. What is the mass-to-charge ratio m/q of the particle in terms of lambda, v, and r?
Answer #1

Electric field due to infinite line charge:

E= λ/2πrε0, where λ=Linear charg density= q/l

Here centripetal force is provided by electric feld

F =q×E =q×λ/2πrε0

F =mv2/r=q×λ/2πrε0

or m/q= q×λ/2πε0×V2 Answer.